OpenStudy (anonymous):

Which number equals the square root of 225? 10 * the square root of 2 15 25 15 * the square root of 5

OpenStudy (anonymous):

Factor to get this solution.\[225=3^25^2 \implies \sqrt{225}=15\]

OpenStudy (anonymous):

thank you

OpenStudy (jiteshmeghwal9):

\[\Huge{\sqrt{225}}\implies15\]

OpenStudy (anonymous):

thank you!

OpenStudy (jiteshmeghwal9):